2- Choose the correct aaswer (Write dowu the correct answer letter AND valuc at your answer a. | 1f random variable Xhas a mea Mr O a standard deviation ơx 4, and (C) 2 and σ r 16 (1) μ,--3 and σ r-8 dE) none of the above t. b. If the continuous random variable X is uniformly distributed witlh a mean of 45 and a variance of 75. The probability that X lies between 40 and 50 is: (C) 2/3 (D) 1/2 (E) none of the above t If a random variable X has a PDE fxCx) 0 elsewhere then EX) i (C) 4/5 (D) 5/4 (E) none of the above p t. d. If a random variable X has a PDF fx(x) 2(x 1) 1sxs2 elsewhere , then σ/s: VAR (E) none of the above e. Let X be a real-valued, continuous raudom variable and let X. Then, If y 0, then the (A) 1/20 ( 208 (D) 1/24 cumulative distribution function Fo) C DP

With explanation please.
Answer #1

a)

\muy =E(2-2X)=2-2E(X)=2-\mux =2-2*0 =2

\sigmay=SD(2-2X)=2*\sigmax=2*4 =8

option E is correct

b)

here as mean =(a+b)/2 =45

a+b=90........(1)

and variance =(b-a)2/12 =75

b-a =30 .........(2)

solving above equations ; a =30 ; b=60

therefore probability X lie between 40 and 50 =(50-40)/(60-30)=10/30=1/3

option A

c)

E(X)=\int_{1}^{2}x f(x) dx= \int_{1}^{2}2(x2-x) dx =2*(x3/3-x2/2)|21 =5/3

option B

d)

E(X2)=\int_{1}^{2}x2 f(x) dx= \int_{1}^{2}2(x3-x2) dx =2*(x4/4-x3/3)|21 =17/6

hence \sigma ^{2}x =E(X2)-(E(X))2 =1/18

option C:

e)

Cumulative funcrtion F(y) =P(Y<y)=P(X2<y) =P(-√y <X<√y) =\int_{-\sqrt{y}}^{\sqrt{y}} fx(x) dx

= Fx(x)|√y-√y   =Fx(√y)-Fx(-√y)

option II